OpenStudy (anonymous):

The table below shows the number of students at Wadesville High School with cell phones, y, over x years. Write an exponential equation that best models these data. Year Cell users 2000 146 2001 183 2002 228 2003 285 2004 356 2005 446 2006 557

ganeshie8 (ganeshie8):

the function will look like below : \(y = 146(B)^{x}\)

ganeshie8 (ganeshie8):

you need to find the base \(B\)

ganeshie8 (ganeshie8):

B = 183/146 ~ 1.25 = 5/4

ganeshie8 (ganeshie8):

does that make some sense ?

OpenStudy (anonymous):

Yes it does and thank you again for the help because I really needed it

ganeshie8 (ganeshie8):

np.. u wlc :)

ganeshie8 (ganeshie8):

so the required exponential function will be : \(y = 146(\frac{5}{4})^{x} \)

ganeshie8 (ganeshie8):

146 is the start value 5/4 is the base and x is the number of years elapsed from 2000